Find the x-intecepts of y=225-x^2

Answer by ewatrrr(24785)


Hi
Finding the x-intercepts of
y= 225-x^2 |when y = 0
0 = 225 -x^2
x^2 = 225
x = ± 15
x-intercepts are (15,0) and (-15,0)